OpenStudy (anonymous):

The region bounded by the given curves is rotated about the specified axis. Find the volume V of the resulting solid by any method. x = (y − 5)^2, x = 4; about y = 3 I am trying to use the disc method but I am not sure how I would.

OpenStudy (anonymous):

Anyone want to help me set up the integral?

OpenStudy (anonymous):

Anyone? :( .

OpenStudy (anonymous):

Well I found the intersection points for y which are y=3 and y=7.

OpenStudy (tkhunny):

Good. This is a nice problem. Have you heard of Pappus's Theorem?

OpenStudy (anonymous):

No.

OpenStudy (tkhunny):

Too bad. It is much simpler with the Centroid. Another lesson on another day, perhaps. The Whole thing: \(\pi\int\limits_{0}^{4}(y_{1}-3)^{2}\;dx\) The Extra thing: \(\pi\int\limits_{0}^{4}(y_{2}-3)^{2}\;dx\) Do you see the nature of the problem that requires two definitions for 'y'? It's NOT a function!

OpenStudy (anonymous):

Yeah, I realized that.

OpenStudy (tkhunny):

Perfect. What are the two definitions?

OpenStudy (anonymous):

Wait you lost me. Never mind.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

\(x = (y-5)^{2}\) Solve for \(y\). You should get TWO results.

OpenStudy (anonymous):

yeah y=y-5 and y=5-y

OpenStudy (anonymous):

Wait. WHat am I doing...

OpenStudy (anonymous):

Braid dead lol.

OpenStudy (tkhunny):

?? Where did the \(x\) go? \(\sqrt{x} = y-5\;for\;y>5\) \(\sqrt{x} = 5-y\;for\;y<5\) Keep solving!

OpenStudy (anonymous):